Find two positive real numbers whose sum is 26 and product is maximum.
Answer by ikleyn(52792)

.
Find two positive real numbers whose sum is 26 and product is maximum.
~~~~~~~~~~~~~~~~


This pair of two positive real numbers is (13,13).


For any other pair of real numbers with the sum of 26,  these numbers are of the form  (13+x,13-x),

and their product  (13+x)*(13-x) = 169-x^2  is LESS THAN 13*13= 169.

Solved, answered and explained.